OpenStudy (anonymous):

whats the equation to find the indicated value? a1=-3 and d=5, determine a15

jimthompson5910 (jim_thompson5910):

If you're dealing with an arithmetic sequence (which it looks like you are), then you use \[\Large a_{n} = a_{1} + d(n-1)\]

jimthompson5910 (jim_thompson5910):

a1 = -3 and d = 5 \[\Large a_{n} = a_{1} + d(n-1)\] \[\Large a_{n} = -3 + 5(n-1)\] then plug in n = 15 \[\Large a_{15} = -3 + 5(15-1)\] \[\Large a_{15} = 67\]

OpenStudy (anonymous):

would \[an=a1+(n-1)d \] be the same equation cuz thats one of the equations my teacher gave me but i wasnt sure?

jimthompson5910 (jim_thompson5910):

yes it is (n-1)d is saying "(n-1) times d" and d(n-1) is the same thing just in a different order

jimthompson5910 (jim_thompson5910):

this is true because you can multiply two numbers in any order ex: 3*2 = 2*3 = 6

OpenStudy (anonymous):

alright thank you!

jimthompson5910 (jim_thompson5910):

np